# Cross-lingual Offensive Language Detection: A Systematic Review of Dataset, Approach and Challenge

### v2

@@ 2.1. Cross-linguality in Offensive Language Detection
Task Definition and Workflow. The task focuses on the identification and categorisation of offensive or hateful text across different languages. The primary objective for CLTL is to leverage the knowledge acquired from one language (source language with more resources) to enhance the HS detection in another language (target language with less resources), especially when the labelled data in the target language is scarce or unavailable (Pamungkas et al., 2021b). The task workflow can be thought of as a four-stage process:
@@ 4. Multilingual Hate Speech Datasets
Data Sources. Most studies collect data from accessible social media such as Twitter and Facebook. Our review reveals a pronounced reliance on popular social media. Twitter encompasses 47.19% of the datasets, followed by Youtube and Facebook (11.24% and 10.11% respectively). While these mainstream platforms dominate, news websites like Fox News and NAVER news, and open forums like Reddit and 2ch, also contribute to the multilingual datasets. And there is a rich diversity in the data source origins. Sources like Weibo (Chinese), 2ch (Russian), g1.globo (Brazilian Portuguese) and Eesti Ekspress (Estonian) represent non-English speaking regions as more culturally specific platforms. Some datasets originate from platforms known for controversial content or specific user bases, such as Stormfront and alt-right websites.
@@
EXIST@IberLEF (Rodr´ıguez-Sánchez et al., 2021) Sexism Identification in Social Networks 2021 sexism en, es 31
@@
EXIST@IberLEF (Rodr´ıguez-Sánchez et al., 2022) Sexism Identification in Social Networks 2022 sexism en, es 19
@@
Prominent topics include offense and HS. Specifically, GermEval competitions in 2018 (Wiegand et al., 2018) and 2019 (Strus̈ et al., 2019) centre around identifying offensive content in German tweets. OffensEval, initially in 2019 focused only on English content (Zampieri et al., 2019), expanded to include multiple languages in the subsequent edition in 2020 (Zampieri et al., 2020). The task in OSACT4 (Mubarak et al., 2020), the 4th Workshop on Open-Source Arabic Corpora and Processing Tools, focuses on offensive language detection in Arabic, especially its dialectal forms. HaSpeeDe has consistently addressed HS with various general and particular topics (Bosco et al., [n. d.]; Sanguinetti et al., 2020; Lai et al., 2023). For instance, HaSpeeDe (Bosco et al., [n. d.]) and HaSpeeDe 2 (Sanguinetti et al., 2020) identify HS against immigrants and Muslims, while HaSpeeDe 3 (Lai et al., 2023) explores HS in strongly polarised debates, particularly political and religious topics. Similarly, the HASOC series, recurring annually from 2019 to 2023, identify HS and offensive content across various languages (Mandl et al., 2019, 2021; Modha et al., 2021; Satapara et al., 2023). TRAC workshops, held in 2018 (Kumar et al., 2018) and 2020 (Kumar et al., 2020), tackled trolling, aggression, and cyberbullying. Other competitions, like PolEval (Ptaszynski et al., 2019) and UrduThreat (Amjad et al., 2022), focus on specific languages (such as Polish and Urdu) or topics (e.g., cyberbullying and threatening languages). The Jigsaw Multilingual Toxic Comment Classification competition on Kaggle attracted over 1600 participants to build multilingual models using English-only training data. Another Kaggle competition, IIIT-D Multilingual Abusive Comment Identification, focuses on abusive comments in Indic languages. Additionally, some tasks focus on more specific topics, such as sexism and misogyny. AMI is organised by IberEval (Fersini et al., 2018b) and Evalita (Fersini et al., 2018a) in 2018, specifically addressing Spanish and Italian misogynous content respectively. EXIST is held for three consecutive years from 2021 to 2023 and aims to identify sexism in social networks (Rodr´ıguez-Sánchez et al., 2021, 2022; Plaza et al., 2023). Most of these benchmark datasets originate from social media platforms, and often conduct annotations that go beyond simple binary labels to finer-grained ones.
@@ 5. Cross-lingual Resources
Machine Translation and Transliteration Tools. Translation refers to converting content from one language into another, whereas transliteration is the process of converting text from one script into another, in both cases preserving the meaning of the original text (Hande et al., 2021b). Machine translation tools can translate datasets from one language to another, such as Google Translate, [^6] https://translate.google.com Microsoft Translator, [^7] https://translator.microsoft.com/ DeepL [^8] https://www.deepl.com/en/translator) and machine translation models (e.g., mBART (Liu et al., 2020) and mT5 (Xue et al., 2021)). Besides, machine transliteration tools are essential for code-mixed datasets, such as Google Transliteration, [^9] https://www.google.co.in/inputtools/try/ Microsoft Transliteration, [^10] https://www.microsoft.com/en-us/translator/business/translator-api/ AI4Bharat transliteration application, [^11] https://github.com/AI4Bharat/IndianNLP-Transliteration and transliteration python packages (e.g., indic-transliteration [^12] https://github.com/indic-transliteration/indic_transliteration_py). However, in offensive language detection, some translation tools may lead to translation errors or cultural nuances that are inconsistent with the original meaning (Pamungkas et al., 2021a). For instance, derogatory terms or slurs without direct equivalents in other languages, or their severity might differ across cultures. Additionally, idiomatic expressions that convey abuse in one language might lose their offensive meaning when translated literally. Therefore, translation quality is crucial to maintain the effectiveness and robustness of cross-lingual models.
@@ 6.1. Instance Transfer
Pairs of instances in source and target languages that share identical labels are referred to as corresponding instances. All data-based transfer techniques necessitate the presence of corresponding texts/labels or a mechanism to generate them. Techniques for instance transfer are introduced distinctly for both text and label levels. For label-based transfer, the primary strategies include annotation projection and pseudo-labelling. Meanwhile, For text-based transfer, the predominant approaches are machine translation and text alignment.
@@ 6.1.1. Label-Based Transfer
It is a semi-supervised learning technique, where a model trained on labelled source data is used to make predictions on unlabeled target data (Lee et al., 2013). High-confidence predictions are treated as “pseudo-labels” for the unlabelled target data. By using high-confidence predictions from a model trained in the source language, it is able to generate labelled target data in another language. This augmented target data can then be used to train or fine-tune models for offensive language detection in the target language. To generate pseudo labels, some studies use an ensemble-based approach to generate labels based on majority voting for unlabelled target datasets, and use bootstrapped target dadasets to further fine-tune the trained model with source training samples (Bigoulaeva et al., 2021, 2022, 2023). Hande et al. (2021b) directly use a pre-trained multilingual language model (i.e. XLM-R (Conneau et al., 2020)) to predict the pseudo labels. Zia et al. (2022) firstly fine-tune a pre-trained XLM-R on gold-labeled source language data, and then use it to create a new psuedo-labelled dataset in the target language.
@@ 6.3.1. Transfer Scenarios
Scenario 1: Zero-Shot Transfer. It is a promising scenario for cross-lingual offensive language detection when labelled data in the target language is scarce or lacking, by enabling detection in a language not seen during training. The model is trained on labelled source data from one or more languages not including the target language, and is then tested on the unseen target language. This is sometimes called direct or model transfer. We found 25 papers under this transfer paradigm, including subtypes:
@@ 7.1.1. Low-resource and Underrepresented Languages
A central challenge in cross-lingual offensive language detection is the limited availability of resources for many non-mainstream languages. As illustrated in Figure 7, the majority of existing studies focus on English as the source language and transfer primarily to typologically similar European languages such as Spanish, German, Italian, and French. In contrast, languages from other families (particularly African, Indigenous, and smaller Asian languages) remain severely underrepresented in both datasets and modelling efforts. This imbalance reflects broader trends in multilingual PLMs, where training corpora disproportionately favour high-resource languages (Nozza, 2021; Röttger et al., 2022a). As a result, CLTL systems often exhibit strong performance for well-represented languages but much weaker transferability for low-resource or distant languages (Deshpande et al., 2022; Eronen et al., 2022). The scarcity of labelled data, the uneven distribution of languages in PLMs, and the difficulty of transferring knowledge across typologically distant families represent major obstacles.
@@ 7.3.1. Limited Ability of Multilingual Pre-Trained Language Models
Cross-lingual models, especially multilingual PLMs, often face challenges related to their generalisation abilities and model bias (Pamungkas et al., 2021a; Hahn et al., 2021; Pelicon et al., 2021b). The generalisability of transformer-based multilingual PLMs can be inconsistent, especially for typologically diverse languages, because they might be pre-trained with a highly focused set of languages or some languages with insufficient training samples (Deshpande et al., 2022). This can lead to bias and instability in model performance, depending on the model architectures, topical focus, linguistic distance, or cultural context in datasets (Mozafari et al., 2022; Pamungkas et al., 2021a). For example, PLMs may exhibit systematic bias when handling dialects, minority languages with various linguistic distances, and content towards marginalised communities. Since such varieties could be under-represented in pre-training corpora, PLMs tend towards standardised forms of dominant languages. This can result in degraded model performance, erasure of dialect-specific expressions, or misinterpretation of community-specific discourse. Model overfitting across target languages is another important issue. Although multilingual models are less prone to overfitting on dataset-specific features than monolingual models, they will achieve poorer performance than monolingual ones in higher-resource settings, and they may require very different calibrations and adaptations across languages (Röttger et al., 2022a, b).
@@ 7.4.2. Data Annotation
Current evidence suggests that incremental and selective annotation works well: starting with small subsets and iteratively expanding the dataset increases efficiency while avoiding redundant labelling (Röttger et al., 2022a). In contrast, large-scale undirected annotation does not work well, as benefits quickly diminish relative to cost. To advance annotation quality, researchers should recruit diverse annotators and provide culturally contextualised guidelines, which helps mitigate annotation bias. Moreover, active learning and semi-supervised strategies have been shown to optimise annotator effort by focusing on ambiguous or uncertain cases (Röttger et al., 2022a). Expert feedback is particularly valuable for subtle and implicit abuse (Vitiugin et al., 2021), but annotator well-being and ethical safeguards must also be prioritised, providing appropriate training and support to avoid potential harm to them.
@@ 7.4.4. Multilingual Pre-Trained Language Models
Multilingual PLMs can be applied to low-resource target languages without further training, thus bridging the data availability gap between high- and low-resource languages. However, there are significant computational costs during training and fine-tuning due to their large number of parameters (Pikuliak et al., 2021). Future research is geared towards optimising these multilingual PLMs for enhanced efficiency and interpretability. Additionally, two strategies are emerging to improve the generalisability and scalability of multilingual PLMs so that they can scale to handle large datasets or multiple languages simultaneously (Arango et al., 2022). The first emphasises pre-training models using data from relevant sources or language communities (such as social platforms and hateful domains), as demonstrated by XLM-T (Montariol et al., 2022) and AbuseXLMR (Gupta et al., 2022). This can mitigate model behaviour bias across different language communities. The second strategy focuses on models based on specific low-resource languages, dialects or those from similar language families, as seen with MuRIL for Indic languages (Das et al., 2022) and AraBERT for dialects (Husain and Uzuner, 2022), which requires intentional inclusion of dialectal and under-represented language data in both training and evaluation. Meanwhile, it is important to evaluate transfer effectiveness across both similar and distant language pairs, rather than relying solely on high-resource or typologically similar examples. Otherwise, multilingual PLMs risk reinforcing existing performance gaps, performing well for Indo-European languages while neglecting typologically distant or underrepresented languages.
@@ 7.4.5. Cross-lingual Training Strategies
Various cross-lingual training strategies have been proven effective. Adaptive training techniques can dynamically adjust to the specificities of different languages and dialects, as well as bridge the gap between offensive language detection and PLMs. By leveraging meta-learning or few-shot learning, models can be trained to quickly adapt to new low-resource languages with few labelled data (Mozafari et al., 2022). Furthermore, collaborative multi-task learning can make models more generalisable in cultural variations and resistant to overfitting by training them and sharing representations on multiple auxiliary tasks simultaneously across different languages or similar NLP tasks (Montariol et al., 2022). The use of adversarial training, where models are trained against adversarial examples, can also enhance their robustness and generalisation across languages. This is especially important for offensive language, as attackers often use subtle language tricks to obfuscate and bypass detection (Shi et al., 2022; Jin et al., 2020; Ryzhova et al., 2022). Moreover, cascade learning can greatly bolster model performance and enhance out-of-domain generalisability by initially training models on source data or more diverse datasets and subsequently fine-tuning the results by simultaneously or progressively incorporating target language data (Röttger et al., 2022a; Pelicon et al., 2021b). Besides, Vitiugin et al. (2021) highlight that human feedback can be of great value to help detect subtleties of abuse and phrases during model training, Ranasinghe and Zampieri (2021c) explore language-specific preprocessing like segmentation in morphologically rich languages such as Arabic and Turkish, and multi-source training, with augmented datasets, is also able to capture the diversities and commonalities of linguistic patterns across languages (Hande et al., 2021a; Markov et al., 2021; Chiril et al., 2019). Future research should focus on systematically comparing these strategies and identifying which are most effective under varying resource and language-family conditions.

### v3

@@ 2.1. Cross-linguality in Offensive Language Detection
Task Definition and Workflow. The task focuses on identifying and categorising offensive or hateful text across languages. The objective of CLTL is to leverage the knowledge acquired from one language (source language with more resources) to enhance HS detection in another language (target language with less resources), especially when the labelled data in the target language is scarce or lacking (Pamungkas et al., 2021b). The task workflow can be thought of as a four-stage process:
@@ 4. Multilingual Hate Speech Datasets
Data Sources. Most studies collect data from social media such as Twitter (now X) and Facebook. Twitter encompasses 47.19% of the datasets, followed by Youtube (11.24%) and Facebook (10.11%). While these mainstream platforms dominate, news websites like Fox News and NAVER news, and open forums like Reddit and 2ch, also contribute to multilingual datasets. Sources like Weibo (Chinese), 2ch (Russian), g1.globo (Brazilian Portuguese) and Eesti Ekspress (Estonian) represent non-English speaking regions as more culturally specific platforms. Some datasets originate from platforms known for controversial content or specific user bases, such as Stormfront and alt-right websites.
@@
EXIST@IberLEF (Rodríguez-Sánchez et al., 2021) Sexism Identification in Social Networks 2021 sexism en, es 31
@@
EXIST@IberLEF (Rodríguez-Sánchez et al., 2022) Sexism Identification in Social Networks 2022 sexism en, es 19
@@
Prominent topics include offense and HS. Specifically, GermEval competitions in 2018 (Wiegand et al., 2018) and 2019 (Strus̈ et al., 2019) centre around identifying offensive content in German tweets. OffensEval, initially in 2019 focused only on English content (Zampieri et al., 2019), expanded to include multiple languages in the subsequent edition in 2020 (Zampieri et al., 2020). The task in OSACT4 (Mubarak et al., 2020), the 4th Workshop on Open-Source Arabic Corpora and Processing Tools, focuses on offensive language detection in Arabic, especially its dialectal forms. HaSpeeDe has consistently addressed HS with various general and particular topics (Bosco et al., [n. d.]; Sanguinetti et al., 2020; Lai et al., 2023). For instance, HaSpeeDe (Bosco et al., [n. d.]) and HaSpeeDe 2 (Sanguinetti et al., 2020) identify HS against immigrants and Muslims, while HaSpeeDe 3 (Lai et al., 2023) explores HS in strongly polarised debates, particularly political and religious topics. Similarly, the HASOC series, recurring annually from 2019 to 2023, identify HS and offensive content across various languages (Mandl et al., 2019, 2021; Modha et al., 2021; Satapara et al., 2023). TRAC workshops, held in 2018 (Kumar et al., 2018) and 2020 (Kumar et al., 2020), tackled trolling, aggression, and cyberbullying. Other competitions, like PolEval (Ptaszynski et al., 2019) and UrduThreat (Amjad et al., 2022), focus on specific languages (such as Polish and Urdu) or topics (e.g., cyberbullying and threatening languages). The Jigsaw Multilingual Toxic Comment Classification competition on Kaggle attracted over 1600 participants to build multilingual models using English-only training data. Another Kaggle competition, IIIT-D Multilingual Abusive Comment Identification, focuses on abusive comments in Indic languages. Additionally, some tasks focus on more specific topics, such as sexism and misogyny. AMI is organised by IberEval (Fersini et al., 2018b) and Evalita (Fersini et al., 2018a) in 2018, specifically addressing Spanish and Italian misogynous content respectively. EXIST is held for three consecutive years from 2021 to 2023 and aims to identify sexism in social networks (Rodríguez-Sánchez et al., 2021, 2022; Plaza et al., 2023). Most of these benchmark datasets originate from social media platforms, and often conduct annotations that go beyond simple binary labels to finer-grained ones.
@@ 5. Cross-lingual Resources
Machine Translation and Transliteration Tools. Translation refers to converting content from one language into another while preserving meaning of the original text, whereas transliteration is the process of converting text from one script into another by preserving phonetic form rather than semantic content (Hande et al., 2021b; Mathur et al., 2018b). Machine translation tools can translate datasets from one language to another, such as Google Translate, [^6] https://translate.google.com Microsoft Translator, [^7] https://translator.microsoft.com/ DeepL [^8] https://www.deepl.com/en/translator) and machine translation models (e.g., mBART (Liu et al., 2020) and mT5 (Xue et al., 2021)). Besides, machine transliteration tools are essential for code-mixed datasets (Kumar and Sachdeva, 2022; Mathur et al., 2018b; Hande et al., 2021b), such as Google Transliteration, [^9] https://www.google.co.in/inputtools/try/ Microsoft Transliteration, [^10] https://www.microsoft.com/en-us/translator/business/translator-api/ AI4Bharat transliteration application, [^11] https://github.com/AI4Bharat/IndianNLP-Transliteration and transliteration python packages (e.g., indic-transliteration [^12] https://github.com/indic-transliteration/indic_transliteration_py). However, in offensive language detection, some translation tools may lead to translation errors or cultural nuances that are inconsistent with the original meaning (Pamungkas et al., 2021a). For instance, derogatory terms or slurs without direct equivalents in other languages, or their severity might differ across cultures. Additionally, idiomatic expressions that convey abuse in one language might lose their offensive meaning when translated literally. Therefore, translation quality is crucial to maintain the effectiveness and robustness of cross-lingual models.
@@ 6.1. Instance Transfer
Pairs of instances in source and target languages that share identical labels are referred to as corresponding instances. All data-based transfer techniques need corresponding texts/labels or a mechanism to generate them. Techniques for instance transfer are introduced distinctly at the text and label levels. For label-based transfer, labels from a source language are projected onto target-language data (e.g., via annotation projection or pseudo labelling), without transferring the source text. In contrast, text-based transfer is to transfer or generate text across languages (e.g., via machine translation or text alignment), preserving the original labels and creating new training instances in the target language. These two strategies differ in which component of an instance is projected across languages, but both establish correspondence between source and target data.
@@ 6.1.1. Label-Based Transfer
It is a semi-supervised learning technique, where a model trained on labelled source data is used to make predictions on unlabeled target data (Lee et al., 2013). High-confidence predictions, namely predicted labels with high probability scores, are treated as “pseudo-labels” for the unlabelled target data. By using high-confidence predictions from a model trained in the source language, it is able to generate labelled target data in another language. This augmented target data can then be used to train or fine-tune models for offensive language detection in the target language. To generate pseudo labels, some studies use an ensemble-based approach to generate labels based on majority voting for unlabelled target datasets, and use bootstrapped target dadasets to further fine-tune the trained model with source training samples (Bigoulaeva et al., 2021, 2022, 2023). Hande et al. (2021b) directly use a pre-trained multilingual language model (i.e. XLM-R (Conneau et al., 2020)) to predict the pseudo labels. Zia et al. (2022) firstly fine-tune a pre-trained XLM-R on gold-labeled source language data, and then use it to create a new psuedo-labelled dataset in the target language.
@@ 6.3.1. Transfer Scenarios
Scenario 1: Zero-Shot Transfer. It is a promising scenario when labelled data in the target language is scarce or lacking, by enabling detection in a language not seen during training. The model is trained on labelled source data from one or more languages not including the target language, and is then tested on the unseen target language. This is sometimes called direct or model transfer. We found 25 papers under this transfer paradigm, including subtypes:
@@ 7.1.1. Low-resource and Underrepresented Languages
A central challenge is the limited availability of resources for many non-mainstream languages. As illustrated in Figure 7, the majority of studies focus on English as the source language and transfer primarily to typologically similar European languages (e.g., Spanish, German, French). In contrast, languages from other families (particularly African, Indigenous, and smaller Asian languages) remain underrepresented in both datasets and modelling efforts. This imbalance reflects broader trends in multilingual PLMs, where training corpora disproportionately favour high-resource languages (Nozza, 2021; Röttger et al., 2022a). As a result, CLTL systems often exhibit strong performance for well-represented languages but much weaker transferability for low-resource or distant languages (Deshpande et al., 2022; Eronen et al., 2022). The scarcity of labelled data, the uneven distribution of languages in PLMs, and the difficulty of transferring knowledge across typologically distant families represent major obstacles.
@@ 7.3.1. Limited Ability of Multilingual Pre-Trained Language Models
Cross-lingual models, especially multilingual PLMs, often lack generalisability and exhibit model bias (Pamungkas et al., 2021a; Hahn et al., 2021; Pelicon et al., 2021b). The generalisability of transformer-based multilingual PLMs can be inconsistent, especially for typologically diverse languages, because they might be pre-trained with a highly focused set of languages or some languages with insufficient training samples (Deshpande et al., 2022). This can lead to bias and instability in model performance, depending on the model architectures, topical focus, linguistic distance, or cultural context in datasets (Mozafari et al., 2022; Pamungkas et al., 2021a). For example, PLMs may exhibit systematic bias when handling dialects, minority languages with various linguistic distances, and content towards marginalised communities. Since such varieties could be under-represented in pre-training corpora, PLMs tend towards standardised forms of dominant languages. This can result in degraded model performance, erasure of dialect-specific expressions, or misinterpretation of community-specific discourse. Model overfitting across target languages is another important issue. Although multilingual models are less prone to overfitting on dataset-specific features than monolingual models, they will achieve poorer performance than monolingual ones in higher-resource settings, and they may require very different calibrations and adaptations across languages (Röttger et al., 2022a, b).
@@ 7.4.2. Data Annotation
Research suggests that incremental and selective annotation works well: iteratively expanding small subsets increases efficiency while avoiding redundant labelling (Röttger et al., 2022a). In contrast, large-scale undirected annotation does not work well, as benefits quickly diminish relative to cost. To advance annotation quality, researchers should recruit diverse annotators and provide culturally contextualised guidelines, which mitigates annotation bias. Moreover, active learning and semi-supervised strategies have been shown to optimise annotator effort by focusing on ambiguous or uncertain cases (Röttger et al., 2022a). Expert feedback is particularly valuable for subtle and implicit abuse (Vitiugin et al., 2021), but annotator well-being and ethical safeguards must also be prioritised, providing appropriate training and support to avoid potential harm.
@@ 7.4.4. Multilingual Pre-Trained Language Models
Multilingual PLMs can be applied to low-resource target languages without further training, thus bridging the data availability gap between high- and low-resource languages. However, their large number of parameters carries significant computational costs during training and fine-tuning (Pikuliak et al., 2021). Future research is geared towards optimising these multilingual PLMs for enhanced efficiency and interpretability. Additionally, two strategies are emerging to improve the generalisability and scalability of multilingual PLMs so that they can scale to handle large datasets or multiple languages simultaneously (Arango et al., 2022). The first emphasises pre-training models using data from relevant sources or language communities (such as social platforms and hateful domains), as demonstrated by XLM-T (Montariol et al., 2022) and AbuseXLMR (Gupta et al., 2022). This can mitigate model bias across language communities. The second strategy focuses on models based on specific low-resource languages, dialects or those from similar language families, as seen with MuRIL for Indic languages (Das et al., 2022) and AraBERT for dialects (Husain and Uzuner, 2022), which requires inclusion of dialectal and under-represented language data in both training and evaluation. Meanwhile, it is important to evaluate transfer effectiveness across both similar and distant language pairs, rather than relying solely on high-resource or typologically similar examples. Otherwise, multilingual PLMs risk reinforcing existing performance gaps, performing well for Indo-European languages while neglecting typologically distant or underrepresented languages.
@@ 7.4.5. Cross-lingual Training Strategies
Various cross-lingual training strategies have been proven effective. Adaptive training techniques can dynamically adjust to the specificities of languages and dialects, as well as bridge the gap between offensive language detection and PLMs. By leveraging meta-learning or few-shot learning, models can be trained to quickly adapt to new low-resource languages with limited labelled data (Mozafari et al., 2022). Furthermore, collaborative multi-task learning can make models more generalisable in cultural variations and resistant to overfitting by training them and sharing representations on multiple auxiliary tasks simultaneously across languages or similar NLP tasks (Montariol et al., 2022). The use of adversarial training, where models are trained against adversarial examples, can also enhance their robustness and generalisation. This is especially important for offensive language, as attackers often use subtle language tricks to obfuscate and bypass detection (Shi et al., 2022; Jin et al., 2020; Ryzhova et al., 2022). Moreover, cascade learning can greatly bolster model performance and enhance out-of-domain generalisability by initially training models on source data or more diverse datasets and subsequently fine-tuning the results by simultaneously or progressively incorporating target language data (Röttger et al., 2022a; Pelicon et al., 2021b). Besides, Vitiugin et al. (2021) highlight that human feedback can be of great value to help detect subtleties of abuse and phrases during model training, Ranasinghe and Zampieri (2021c) explore language-specific preprocessing like segmentation in morphologically rich languages (e.g., Arabic, Turkish), and multi-source training, with augmented datasets, is also able to capture the diversities and commonalities of different linguistic patterns (Hande et al., 2021a; Markov et al., 2021; Chiril et al., 2019). Future research should focus on systematically comparing these strategies and identifying which are most effective under varying resource and language-family conditions.